OpenStudy (anonymous):

what basic trigonometric identity would you use to verify that tan x cos x= sin x? a. cos x=1/sec x b. sin s=1/csc x c. cos^2x + sin^2x=1 d. tan x = sin x/ cos x

OpenStudy (anonymous):

D

OpenStudy (jziggy):

Converting everything to sines and cosines will usually work. So, tan(x)cos(x) = sin(x) (sin(x)/cos(x))cos(x) = sin(x) sin(x) = sin(x)